# Self-Training Meets Consistency: Improving LLMs’ Reasoning With Consistency-Driven Rationale Evaluation

### v3

@@ Appendix E Prompts
Figure 7: Prompt template for rationale generation and inference.
[Instruction]
@@
Figure 8: Prompt template for deriving an answer prediction a given rationale.
[Instruction]
@@
(Rationale here)
@@
Figure 9: Prompt template for follow-up questions.
E.1 Rationale Generation and Evaluation
@@ E.2 Follow-up questions
E.3 Prompt and Example of Qualitative Analysis with FLASK
@@
Please give feedback on the assistant’s responses. Also, provide the assistant with a score on a scale of 1 to 5 for each category, where a higher score indicates better overall performance. Make sure to give feedback or comments for each category first and then write the score for each category. Only write the feedback corresponding to the score rubric for each category. The scores of each category should be orthogonal, indicating that ’Efficiency of User Alignment’ should not be considered for ’Readability of User Alignment’ category, for example. Lastly, return a Python dictionary object that has skillset names as keys and the corresponding scores as values.
@@
Appendix F Implementation Details
We use lora rank=16, alpha=16 and target modules = {gate_proj, down_proj, up_proj, q_proj, k_proj, v_proj, o_proj}. We use cosine scheduler and adamW optimizer Loshchilov and Hutter (2017). For saving memory, we use half-precision (fp16) in training $\textbf{{M}}_{\textbf{SFT}}$. During inference, if the model fails to fully generate the answer label within the designated generation length, we clarify the prediction by appending ‘[Answer] Therefore, the answer is’ to the end of the initial output and conduct an additional query. We select models that show the highest performance on the validation set without early-stopping. For Llama 3 8B experiments on ReClor, the best-found hyperparameter values for the supervised fine-tuning stage were: learning rate=8e-4, batch size=8, tolerance=3. For the preference learning stage, the best-found hyperparameter values were $\lambda$=0.6, learning rate=6e-6, max number of steps=5000. Our hardware setting is Intel(R) Xeon(R) Gold 5218R CPU @ 2.10GHz (CPU), and NVIDIA RTX A6000 (GPU). We use vllm Kwon et al. (2023) library for efficient rationale generation and evaluation. We use trl von Werra et al. (2020) library for supervised fine-tuning and preference learning stages. For the datasets we use in this paper, CSQA is under MIT license, and ARC is under CC BY-SA 4.0 license. You can see terms for use of ReClor in here. We use these datasets and the models solely for research purposes.
Appendix G Computational Costs
@@
Appendix H Adjustments in Implementation of Baseline Models

### v4

@@ Appendix F Prompts
  Output  (generated rationale here)
[Answer]
Therefore, the answer is (answer label here).
Figure 7: Prompt template for rationale generation and inference. This template is used for generating rationales and evaluating models in self-training approaches (Table 1), as well as Zero-shot-CoT and Few-shot-CoT models (Table 4).
Input  [Instruction]
@@
  Output  (answer label here).
Figure 8: Prompt template for deriving an answer prediction from a given rationale. The answer prediction is compared to the ground truth to evaluate each generated rationale and calculate the reward $z$ for it.
Input  [Instruction]
@@
Let’s think step by step.
(generated rationale here)
@@
  Output  (the/not the) correct answer.
Figure 9: Prompt template for evaluation using follow-up questions. This template evaluates a given rationale by prompting models to solve a follow-up question based on the rationale. As shown in the input part, the follow-up question asks whether the target option is the correct answer to the original question. Results for all target answer choices are aggregated to validate the given rationale and compute the reward $\tilde{z}$.
Input  [Instruction]
(instruction here)
[Question]
<Passage> (passage here)
<Question> (question here)
Answer Choices:
A. (option A here)
B. (option B here)
C. (option C here)
D. (option D here)
[Answer]
The correct answer is
Output  (answer label here).
Figure 10: Prompt template for direct answer prediction. This template is used to evaluate Zero-shot, Few-shot, and Direct fine-tuning approaches (Table 1). Unlike other templates, it does not require models to generate or utilize rationales.
F.1 Rationale Generation and Evaluation
@@ F.2 Follow-up Questions
[…]
F.4 Evaluating Models
Figure 7 shows the prompt template used for evaluating models in self-training approaches (Table 1) as well as Zero-shot-CoT and Few-shot-CoT models (Table 4). Figure 10 shows the prompt template for direct answering, where models are provided with a question and tasked with predicting the answer directly, without generating rationales. This template is used to evaluate Zero-shot, Few-shot, and direct fine-tuning methods, as detailed in Table 1.
F.5 Prompt and Example of Qualitative Analysis with FLASK
@@
Please give feedback on the assistant’s responses. Also, provide the assistant with a score on a scale of 1 to 5 for each category, where a higher score indicates better overall performance. Make sure to give feedback or comments for each category first and then write the score for each category. Only write the feedback corresponding to the score rubric for each category. The scores of each category should be orthogonal, indicating that ‘Efficiency of User Alignment’ should not be considered for ‘Readability of User Alignment’ category, for example. Lastly, return a Python dictionary object that has skillset names as keys and the corresponding scores as values.
@@
Appendix G Implementation Details
We use lora rank=16, alpha=16 and target modules = {gate_proj, down_proj, up_proj, q_proj, k_proj, v_proj, o_proj}. We use cosine scheduler and adamW optimizer Loshchilov and Hutter (2017). For saving memory, we use half-precision (fp16) in training $\textbf{{M}}_{\textbf{SFT}}$. During inference, if the model fails to fully generate the answer label within the designated generation length, we clarify the prediction by appending ‘[Answer] Therefore, the answer is’ to the end of the initial output and conduct an additional query. We select models that show the highest performance on the validation set without early-stopping. For Llama 3 8B experiments on ReClor, the best-found hyperparameter values for the supervised fine-tuning stage were: learning rate=8e-4, batch size=8, tolerance=3. For the preference learning stage, the best-found hyperparameter values were $\lambda$=0.6, learning rate=6e-6, and max number of steps=5000. Our hardware setting is Intel(R) Xeon(R) Gold 5218R CPU @ 2.10GHz (CPU), and NVIDIA RTX A6000 (GPU). We use vllm Kwon et al. (2023) library for efficient rationale generation and evaluation. We use trl von Werra et al. (2020) library for supervised fine-tuning and preference learning stages. For the datasets we use in this paper, CSQA is under the MIT license, and ARC is under the CC BY-SA 4.0 license. You can see terms for use of ReClor in here. We use these datasets and the models solely for research purposes.
Appendix H Computational Costs
@@
Appendix I Adjustments in Implementation of Baseline Models